OpenStudy (anonymous):

Express in simplest form: 3x + 6y ------- 3x + 3y

OpenStudy (anonymous):

To Undo the bottom part you need to multiply by it opposite sign. 3x+6y 3x-3 ------ * ----- 3x+3 3x-3 Can you try to solve this?

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

Can you explain further? I think I have it from here. Thanks. By "opposite sign" you mean?

jimthompson5910 (jim_thompson5910):

no need to do that, just factor 3 from each expression and cancel

OpenStudy (anonymous):

@Jim, so 3x + 6 y -------- 3x + 3y 3y?

jimthompson5910 (jim_thompson5910):

not quite bryce101

OpenStudy (anonymous):

@jim_thompson5910 That is true but not all problems can be solve "nice and easy" @Bryce101 this should help you. http://www.purplemath.com/modules/radicals4.htm It's a theorem call multiply by its conjugates to make the problem more simple.

jimthompson5910 (jim_thompson5910):

3x + 6 y = 3(x+2y) and 3x+3y = 3(x+y) Does that help?

OpenStudy (anonymous):

Yes thank you both! :)

OpenStudy (anonymous):

@Bryce101 Do you have the Answer? What does it mean to Express in the Simplest form?

jimthompson5910 (jim_thompson5910):

So you should get \[\Large \frac{x+2y}{x+y}\] as your final answer.